OpenStudy (lolacole12):

Complete the equation that is related to the given equation. x – (–6.9) = 42.8 x = 42.8 +

OpenStudy (lolacole12):

@hannabanana04 @hopyhopy

OpenStudy (lolacole12):

@mathmale

OpenStudy (lolacole12):

@just_one_last_goodbye

OpenStudy (lolacole12):

oll give medal @megax

OpenStudy (lolacole12):

@Gabylovesyou

OpenStudy (anonymous):

@lolacole12

OpenStudy (lolacole12):

help

OpenStudy (anonymous):

what do you need help with?

OpenStudy (anonymous):

oh hmmm

OpenStudy (lolacole12):

Complete the equation that is related to the given equation. x – (–6.9) = 42.8 x = 42.8 +

OpenStudy (anonymous):

here X – (–6.9) = 42.8 x + 6.9 = 42.8 x = 35.9

OpenStudy (lolacole12):

so its 35.9

OpenStudy (anonymous):

yep

OpenStudy (lolacole12):

one more

OpenStudy (anonymous):

your not suposed to give direct so i just made it realy easy

OpenStudy (lolacole12):

Which value of x makes the equation true? x – 18.7 = –12 x =

OpenStudy (lolacole12):

i got it though

OpenStudy (lolacole12):

you know this

OpenStudy (anonymous):

oh you got it

OpenStudy (lolacole12):

yes

OpenStudy (anonymous):

ok

OpenStudy (lolacole12):

is it 6.7

OpenStudy (lolacole12):

@megax